One train travels 810 miles in the same amount of time it takes another train to travel 930 miles. The first trains rate is 8 miles per hour slower then the second trains rate, Find the average rate of each train.
Answer by richwmiller(17219)

(r-8)*t=810,
r*t=930
r = 62, t = 15
r-8=54